Based on our understanding of the Aggregate Expenditure (AE) Model (AE), we know that an increase in c1 (where C = c0 + c1Yd) will cause the AE function to become:

Group of answer choices

Flatter and a given change in autonomous consumption (c0) to have a larger effect on output.

Steeper and a given change in autonomous consumption (c0) to have a smaller effect on output.

Flatter and a given change in autonomous consumption (c0) to have a smaller effect on output.

Steeper and a given change in autonomous consumption (c0) to have a larger effect on output.
1
Expert's answer
2021-09-24T11:29:12-0400

Y = c0 + c1Yd + I + G

c= marginal propensity to consume. It captures by how much households increase their consumption if their disposable income increases marginally. It is the slope coefficient and its value can vary from 0 < c1 < 1 .

Y = c0 + c1(Y − T) + I + G

Y = 1/1−c1 [c0 + I + G − c1T]

1/(1-c1) is the multiplier. So as the value of cincreases, it will result in the increasing the value of multiplier and therefore the output Y will have a larger effect. Being the slope coefficient, increase in value of c1  will lead to a steeper shape of the curve.

Hence, the answer is Option D: Steeper and a given change in autonomous consumption (c0) to have a larger effect on output.